During a race on an oval truck, a cars travel at an average speed of 200 km/h. (a) How far did it travel in 45.0 min? (b) Determine its average velocity at the end of its third lap.


1
Expert's answer
2021-03-12T07:13:49-0500

(a)

"d=vt=200\ \dfrac{km}{h}\cdot45\ min\cdot\dfrac{1\ h}{60\ min}=150\ km."

(b) The average velocity of the car is its total displacement divided by the total time taken:


"v_{avg}=\dfrac{Displacement}{Time\ taken}."

Since the car returns to the beginning of the lap its displacement is zero and the average velocity is zero as well.